How can I create an equation or an inequality that can be used for analysis, prediction, or decision making when discussing the weather temperatures from one year to another?
Answer by tutorcecilia(2152)

Take the average temperature of March (or any other month) for three consequative years. Assign "x" to the years and "y" to the temperature. Suppose the temperature is rising each year due to global warming, than the average temperature during March (or any other chosen month), should rise in a linear fashion.
.
For example: The Average Temperature in March was:
x = 1990, 1991, 1992
y = 75......85.......95
.
Than, the points are (0, 75) (1, 85) and (2, 95)
Assuming that this trend in global warming continues indefinately, you can use these points to predict the average temperature in any future year.
.
Find the slope of the line. Pick any two points:
m = (y2-y1)/(x2-x1) [Plug-in any two points and solve for the slope]
m = (85-75)/(1-0) = 5/1=5
.
Use the slope-intercept equation of a line:
y = m x + b [plug-in any values for (x, y). Plug-in the slope]
75 = (5)(0) + b [Solve for "b"]
75 = b [Plug-in the values for (m) and (b)which will not change]
y = 5x+75 [This is the predictor or equation of the line for our proposed global warming in March]
.
y = 5(9)+75 [Pick the year 1999 and plug it in for (x=9)]
y = 120 degrees!!! [Solve to find the average temperature in March of 1999 if global warming continues]